OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 − (x + 2) = 6 − 5(x + 4) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

4-X-2=6-5X-20;then ; 2-X=-14-5X ;then;2+14=-5X+X ;then; 16=-4X ;then "X"=-4

OpenStudy (anonymous):

thnks

hero (hero):

4 - (x+2) = 6 - 5(x+4) Add 5(x+4) to both sides; Subtract 4 from both sides to get: 5(x+4) - (x+2) = 6 - 4 Distribute 5 across (x+4) to get: 5x + 20 -(x+2) = 2 Distribute -1 across(x+2) to get: 5x + 20 - x - 2 = 2 Place like terms together to get 5x - x + 20 - 2 = 2 Combine terms to get 4x + 18 = 2 Subtract 18 from both sides to get 4x = -16 Divide both sides by 4 to get x = -4

OpenStudy (anonymous):

4 − (x + 2) = 6 − 5(x + 4) Hero ...already took over